Silver has 2 commonly occurring isotopes. 51.839% of Silver atoms have 60 neutrons and 47 protons, whereas the other 48.161% has 62 neutrons and 47 protons. Using the isotopic composition provided, and the known masses of protons (1.00727 amu) and neutrons (1.00866 amu), calculate the expected mass of Silver. Round your answer to 3 decimal places.
1
Expert's answer
2020-09-22T03:35:08-0400

1. find the masses of the isotopes:

1-isotope: 60+47=107

2-isotope: 62+47=109

2.now we find the mass of the silver atom. To do this, we multiply the mass of the isotopes by its percentage, add them together, and divide by 100%:

(107*51.839+109*48.161)/100=107.96322≈108

ANSWER: 108